OpenStudy (anonymous):

are there any good tips for taking a calculus 2 exam? What is an easy way to remember how to do trig substitutions?

OpenStudy (amistre64):

... do your homework.

OpenStudy (anonymous):

dont do homework

OpenStudy (jamesj):

Practice, practice, practice.

myininaya (myininaya):

there are only three trig substitution you need if you see something in the form of: x^2-a^2, x^2+a^2, or a^2-x^2. if you see x^2-a^2, use sec(theta)=x/a if you see x^2+a^2 use tan(theta)=x/a if you see a^2-x^2 use sin(theta)=x/a now there are other probably outside of this section different substitution of trig that you might have to use but if just if you see one of those forms you should be able to use what i told you above though sometimes a substitution like this can be avoided when you have something and this form 2x/(x^2-a^2) and you would like to integrate then you just need a normal substitution (or let better phrase an algebraic substitution though a trig substitution would probably turn out okay in the end you are creating more work than there is) sometimes things can get trickier it really is practice and more practice you need but anyways you might do trig substitution elsewhere for example to integrate: \[\int\limits_{}^{}\tan^{-1}(x) dx\] I would let \[x=\tan(\theta)\] \[dx=\sec^2(\theta) d \theta \] so we have \[\int\limits_{}^{}\tan^{-1}(\tan(\theta)) \sec^2(\theta) d \theta=\int\limits_{}^{} \theta \sec^2(\theta) d \theta\] then i would do integration by parts \[\theta \tan(\theta)-\int\limits_{}^{} \tan(\theta) d \theta=\theta \tan(\theta)-\int\limits_{}^{}\frac{\sin(\theta)}{\cos(\theta)} d \theta\] then i would let u=cos(theta) => du=-sin(theta) d theta \[\int\limits_{}^{} \tan^{-1}(x) dx=\int\limits_{}^{}\theta \sec^2(\theta) d \theta=\theta \tan(\theta)-\int\limits_{}^{}\frac{-du}{u}\] \[=\theta \tan(\theta)+\ln|u|+C=\theta \tan(\theta)+\ln|\cos(\theta)|+C\] which we need to back in terms of x using a right triangle recall i let x=tan(theta)  \[=\tan^{-1}(x)+\ln|\frac{adjacent of \theta}{hyp}|+C\] \[=\tan^{-1}(x)+\ln|\frac{1}{\sqrt{x^2+1}}|+C\]

myininaya (myininaya):

oops i forgot reaplacing tan(theta)

myininaya (myininaya):

\[=\tan^{-1}(x) \cdot x+\ln|\frac{1}{\sqrt{x^2+1}}|+C\]

OpenStudy (amistre64):

someones writing themselves a novel ... i hope theres a vampire in it

myininaya (myininaya):

lol